# Probabilistic Forecasting with Coherent Aggregation

### v3

@@ Abstract
Obtaining accurate probabilistic forecasts is an important operational challenge in many applications, like energy management, climate forecast, supply chain planning, and resource allocation. In many of these applications, there is a natural hierarchical structure over the forecasted quantities; and forecasting systems that adhere to this hierarchical structure are said to be coherent. Furthermore, operational planning benefits from accuracy at all levels of the aggregation hierarchy. Building accurate and coherent forecasting systems, however, is challenging: classic multivariate time series tools and neural network methods are still being adapted for this purpose. In this paper, we augment an MQForecaster neural network architecture with a novel deep Gaussian factor forecasting model that achieves coherence by construction, yielding a method we call the Deep Coherent Factor Model Neural Network (DeepCoFactor) model. DeepCoFactor generates samples that can be differentiated with respect to the model parameters, allowing optimization on various sample-based learning objectives that align with the forecasting system’s goals, including quantile loss and the scaled Continuous Ranked Probability Score (CRPS). In a comparison to state-of-the-art coherent forecasting methods, DeepCoFactor achieves significant improvements in scaled CRPS forecast accuracy, with average gains of 15%, as measured on six publicly-available forecasting datasets.
{NoHyper} Equal contribution.{NoHyper} Work done while at Amazon Forecasting Science. Geoffrey was a PhD candidate at UC Berkeley when interning with Amazon.
@@ 1 Introduction
Obtaining accurate forecasts is an important step for long-term planning in complex and uncertain environments, with applications ranging from energy to supply chain management, from transportation to climate prediction [18, 15, 31]. Going beyond point forecasts such as means and medians, probabilistic forecasting provides a key tool for forecasting uncertain future events. This involves, e.g., forecasting that there is a 90% chance of rain on a certain day, or that there is a 99% chance that people will want to buy fewer than 100 items at a certain store on a certain week. Providing more detailed predictions of this form permits finer uncertainty quantification. This in turn permits planners to prepare for different scenarios and to allocate resources depending on their anticipated likelihood and cost structure. This can lead to better resource allocation, improved decision making, and less waste.
In many forecasting applications, there exist natural hierarchies over the quantities one wants to predict, such as energy consumption at various temporal granularities (from monthly to weekly), different geographic levels (from building-level to city-level to state-level), or retail demand for specific items (in a hierarchical product taxonomy). Typically, most or all levels of the hierarchy are important: the bottom levels are key for operational short-term planning, while higher levels of aggregation provide insights into longer-term or broader trends. Moreover, it is often desired that probabilistic forecasts are coherent (or consistent) to ensure efficient decision-making at all levels [18, 24]. Coherence is achieved when the forecast distribution assigns zero probability to forecasts that do not satisfy the hierarchy’s constraints [40, 5, 37] (see Definition 2.1). Designing an accurate model, capable of leveraging information from all hierarchical levels, while enforcing coherence is a well-known and challenging task [20].
Method End-to-End Multivariate Series Inputs Multivariate Output Distribution Arbitrary Learning Objective
@@
The hierarchical forecasting literature has been dominated by two-stage reconciliation approaches, where univariate methods are first fitted and later reconciled towards coherence. For many years, most research focused on mean reconciliation [20, 22, 49, 21, 8, 53, 33]. More recent statistical methods consider coherent probabilistic forecasts through variants of the bootstrap reconciliation technique [5, 40] or the clever use of the Gaussian forecast distributions’ properties [52]. Large-scale applications of hierarchical forecasting require one to simplify over the two-stage reconciliation process by favoring end-to-end approaches that simultaneously fit all levels of the hierarchy, while still achieving coherence. The end-to-end approach refers to training a model constrained to achieve coherence by optimizing directly for accuracy. End-to-end methods offer advantages such as reduced complexity, improved computational efficiency, and enhanced adaptability by streamlining the entire forecasting pipeline into a single, unified model. More importantly, end-to-end models generally achieve better accuracy compared to two-stage models that are first trained independently for optimized accuracy and then made coherent through various reconciliation approaches [43, 37].
@@
To the best of our knowledge, only three methods yield coherent probabilistic forecasts and allow models to be trained in an end-to-end manner: Rangapuram et al. [43], Olivares et al. [37], and Das et al. [9]. (There is also parallel research on hierarchical forecasts with relaxed constraints [17, 41, 25, 26, 48], but we limit our analysis of this line of work since our focus is on strictly coherent forecasting methods.) In particular, Olivares et al. [37] considers a finite mixture of Poisson distributions that captures correlations implicitly through latent variables and does not leverage cross-time series information, and coherence is achieved by constructing probabilistic forecasts for aggregated time series through equality completion [10]. On the other hand, Rangapuram et al. [43] leverages the multivariate time series information, but it achieves coherence through a differentiable projection layer, which could degrade forecast accuracy: it does not directly model correlations between the multivariate outputs, but rather it couples them through its projection layer. Dedicated effort is still necessary to capture these hierarchical relationships to improve forecast accuracy. Such probabilistic methods can benefit from the ability to optimize for arbitrary loss functions through making samples differentiable, as demonstrated by Rangapuram et al. [43]. The capacity to optimize any loss computed from forecast samples can help align the forecasting system’s goals with the neural network’s learning objective.
@@
Summarized, an ideal hierarchical forecasting method should satisfy several desiderata: 1) be end-to-end coherent; 2) model a joint multivariate probability distribution, capturing the intricate relationships between series within the hierarchy; 3) leverage cross time series information to accurately reflect these relationships; and 4) generate differentiable samples, to enable the method to optimize for arbitrary learning objectives that align with the forecasting system’s goals. In this paper, we present a method which satisfies all of these ideal properties; see Table 1 for a summary.
@@
We introduce the Deep Coherent Factor Model Neural Network (DeepCoFactor), a method for producing probabilistic coherent forecasting that satisfies all the desired properties stated above. Our main contributions are the following.
@@
    We design a multivariate Gaussian factor model that achieves forecast coherence exactly by construction, representing the joint forecast distribution over the bottom-level series. This factor model is generic and can be applied to most neural forecasting models with minimal modifications.
@@
    In contrast to other joint distribution models, our factor model samples’ differentiability—enabled by the reparameterization trick—offers greater versatility, allowing to optimize a wider range of learning objectives beyond negative log likelihood. We leverage this versatility by training the model using both the Continuous Ranked Probability Score (CRPS) and the Energy Score [32]. For hierarchical forecasting tasks, aligning the training objective with the CRPS evaluation metric leads to significant improvements in accuracy.
@@
    We combine the multivariate factor model with an augmented MQCNN architecture [51, 36] into the Deep Coherent Factor Model Neural Network (DeepCoFactor). Our augmented MQCNN architecture accepts Vector Autorregresive (VAR) inputs through a cross series multi layer perceptron module.
@@
    DeepCoFactor achieves state-of-the-art performance across six publicly available benchmark datasets, demonstrating a substantial improvement in both probabilistic and mean forecast accuracy. On average, DeepCoFactor outperforms the second-best method by 15 percent in probabilistic forecasting accuracy and delivers a 30 percent increase in mean forecast accuracy.
@@
The aggregation matrix ${\mathbf{A}}_{[a][b]}$ represents the collection of linear transformations for deriving the aggregates, and sums the bottom series to the aggregate levels. The hierarchical aggregation constraints matrix ${\mathbf{S}}_{[i][b]}$ obtained by stacking ${\mathbf{A}}_{[a][b]}$ and the $N_{b}\times N_{b}$ identity matrix ${\mathbf{I}}_{[b][b]}$.
For a simple example, consider $N_{b}=4$ bottom-series, so $[b]=\{1,2,3,4\}$ and $y_{Total,t}=\sum_{i=1}^{4}y_{i,t}$. Figure 1 shows an example of such hierarchical structure, where the multivariate hierarchical time series is defined
@@
The hierarchical forecasting task augments the forecast probability in Eqn. 3 with coherence constraints in  Eqn. 2.1 [6, 40, 36], by restricting the probabilistic forecast space to assign zero probability to non-coherent forecasts. Definition 2.1 formalizes the intuition, and states that the distribution of a given aggregate random variable is exactly the distribution defined as the aggregates of the bottom-series distributions through the summation matrix ${\mathbf{S}}_{[i][b]}$.
@@
Let $(\Omega_{[b]},\mathcal{F}_{[b]},\mathbb{P}_{[b]})$ be a probabilistic forecast space (on the bottom-level series $\mathbf{Y}_{[b]}$), with sample space $(\Omega_{[b]}$, $\mathcal{F}_{[b]})$, and $\mathbb{P}_{[b]}$ a forecast probability. Let $\mathbf{S}_{[i][b]}(\cdot):\Omega_{[b]}\mapsto\Omega_{[i]}$ be the linear transformation implied by the aggregation constraints matrix. A coherent forecast space $(\Omega_{[i]},\mathcal{F}_{[i]},\mathbb{P}_{[i]}$) satisfies
@@
for any set $\mathcal{B}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{F}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}_{[b]}$ and set’s image ${\mathbf{S}}_{[i][b]}(\mathcal{B})\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{F}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}_{[i]}$.
@@
In this work and most hierarchical forecasting literature, the performance of probabilistic forecasts is primarily evaluated by the Continuous Ranked Probability Score (CRPS), e.g. [5, 43, 37, 40, 9, 52]. The CRPS between a target $y$ and distributional forecast $Y$ is defined as
@@
CRPS is commonly used because it is a strictly proper scoring rule [16], and is agnostic of model and distributional assumptions. The metric also gives a summary of peformance on all quantile forecasts [29], as
@@ 3 Methodology
In this section, we describe our main method, the Deep Coherent Factor Model Neural Network (DeepCoFactor) model. It consists of a multivariate probabilistic model, an underlying neural network structure, and an end-to-end model estimation procedure.
3.1 Multivariate Probabilistic Model
Our predicted probabilistic forecasts at all hierarchical levels are jointly represented by a Gaussian factor model. Our neural network maps the known information (past, static and known future) to the location, scale and shared factor parameters, and the forecasted factor model parameters are designed to model correlations between the bottom-level series, while conditioning on all known information. Our factor model [^2] Early work on factor forecast models augmenting neural networks done by Wang et al. [50] does not ensure coherence. combined with the coherent aggregation in Eqn. 10 directly estimates the multivariate probability of bottom-level series ${\mathbf{y}}_{[b][t+1:t+N_{h}]}$ conditioning on historical, known-future, and static covariates ${\mathbf{x}}^{(h)}_{[b][:t]},\;{\mathbf{x}}^{(f)}_{[b][t+1:t+N_{h}]},{\mathbf{x}}^{(s)}_{[b]}$, i.e.,
@@ 3.1 Multivariate Probabilistic Model
After sampling from the multivariate factor we coherently aggregate the clipped outputs of the network,
@@
Any bottom-level multivariate distribution can be transformed into a coherent distribution through coherent aggregation  [^3] Coherent aggregation can be though of a special case of bootstrap reconciliation [40] that only relies on a bottom-level forecast distribution..Given a sample $\mathbf{\hat{y}}_{[b]}\sim\mathbb{P}_{[b]}$, a coherent $\mathbb{P}_{[i]}$ distribution can be constructed with the following sample transformation
@@
In other words, it is enough to aggregate the bottom-level forecasts in a bottom-up manner. We include in Appendix A a proof of the approach’s coherence property, and the covariance structure.
@@ 3.2 Neural Network Architecture
Figure 2: The Deep Coherent Factor Model Neural Network is a Sequence-to-Sequence with Context network that uses dilated temporal convolutions as the primary encoder and multilayer perceptron based decoders for the creation of the multi-step forecast. DeepCoFactor coherently aggregates the samples of the factor model $\mathbf{\tilde{y}}_{[i],\eta,t}=\mathbf{S}_{[i][b]}\mathbf{\hat{y}}_{[b],\eta,t}$. We mark in red the standard normal samples that are parameter-free, the reparametrization trick allows to apply backpropagation through the factor model outputs. DeepCoFactor extends upon the univariate MQCNN, through the cross series multi layer perceptron.
Our framework can adapt to any univariate deep learning forecasting models, so long as they can be made to output factor model parameters. In this paper, we focus on the architecture implementation based on MQCNN [51, 36], because of its outstanding performance in multi-step forecasting problems. Our MQCNN-based architecture has a main encoder which consists of a stack of dilated temporal convolutions, and it is applied to historical information for all series. In addition, it uses a global multi-layer perceptron (MLP) to encode the static and future information. The encoder at time $t$ is described in Eqn. 13 below, and is applied to each disaggregated series:
@@
We use a residual cross series MLP to capture vector autorregressive relationships in the hierarchy with minimal modifications to the architecture:
@@
The DeepCoFactor uses a two-stage MLP decoder: the first decoder summarizes information into the horizon agnostic context $\mathbf{c}^{(ag)}_{[b],t}$ and the horizon specific context $\mathbf{c}^{(sp)}_{[b][h],t}$; and the second stage decoder transforms the contexts into the Factor model parameters $\left(\bm{{\hat{\mu}}}_{[b][h],t},\;\bm{{\hat{\sigma}}}_{[b][h],t},\mathbf{{\hat{F}}}_{[b][k][h],t}\right)$. Eqn. 15 describes the operations:
@@
For the last step, the network composes the factor model samples, using Eqn. 9 and aggregates them (equivalent to bottom-up reconciliation) in Eqn. 10.

### v4

# Coherent Learning Objective Reparameterization

@@ Abstract
Obtaining accurate probabilistic forecasts is an operational challenge in many applications, such as energy management, climate forecasting, supply chain planning, and resource allocation. Many of these applications present a natural hierarchical structure over the forecasted quantities; and forecasting systems that adhere to this hierarchical structure are said to be coherent. Furthermore, operational planning benefits from the accuracy at all levels of the aggregation hierarchy. However, building accurate and coherent forecasting systems is challenging: classic multivariate time series tools and neural network methods are still being adapted for this purpose. In this paper, we augment an MQForecaster neural network architecture with a modified multivariate Gaussian factor model that achieves coherence by construction. The factor model samples can be differentiated with respect to the model parameters, allowing optimization on arbitrary differentiable learning objectives that align with the forecasting system’s goals, including quantile loss and the scaled Continuous Ranked Probability Score (CRPS). We call our method the Coherent Learning Objective Reparametrization Neural Network (CLOVER). In comparison to state-of-the-art coherent forecasting methods, CLOVER achieves significant improvements in scaled CRPS forecast accuracy, with average gains of 15%, as measured on six publicly-available datasets.
{NoHyper} Equal contribution.{NoHyper} Work completed at Amazon Forecasting Science, where Geoffrey interned as a UC Berkeley Ph.D. candidate.
@@ 1 Introduction
Obtaining accurate forecasts is an important step for planning in complex and uncertain environments, with applications ranging from energy to supply chain management, from transportation to climate prediction [19, 16, 32]. Going beyond point forecasts such as means and medians, probabilistic forecasting provides a key tool for predicting uncertain future events. This involves, for example, forecasting that there is a 90% chance of rain on a certain day or that there is a 99% chance that people will want to buy fewer than 100 items in a certain store in a given week. Providing more detailed predictions of this form permits finer uncertainty quantification. This, in turn, allows planners to prepare for different scenarios and allocate resources according to their anticipated likelihood and cost structure. This can lead to better resource allocation, improved decision-making, and less waste.
In many forecasting applications, there exist natural hierarchies over the quantities one wants to predict, such as energy consumption at various temporal granularities (from monthly to weekly), different geographic levels (from building-level to city-level to state-level), or retail demand for specific items (in a hierarchical product taxonomy). Typically, most or all levels of the hierarchy are important: the bottom levels are key for operational short-term planning, while higher levels of aggregation provide insight into longer-term or broader trends. Moreover, probabilistic forecasts are often desired to be coherent (or consistent) to ensure efficient decision-making at all levels [19, 25]. Coherence is achieved when the forecast distribution assigns zero probability to forecasts that do not satisfy the hierarchy’s constraints [6, 41, 38] (see Definition 2.1). Designing an accurate model, capable of leveraging information from all hierarchical levels while enforcing coherence is a well-known and challenging task [21].
Method Coherent End-to-End Multivariate Series Inputs Multivariate Output Distribution Arbitrary Learning Objective
@@
The hierarchical forecasting literature has been dominated by two-stage reconciliation approaches, where univariate methods are first fitted and later reconciled towards coherence. For many years, most research has focused on mean reconciliation [21, 23, 51, 22, 9, 55, 34]. More recent statistical methods consider coherent probabilistic forecasts through variants of the bootstrap reconciliation technique [6, 41] or the clever use of the properties of the Gaussian forecast distributions [54]. A detailed hierarchical forecast review is provided by Athanasopoulos et al. [2]. Large-scale applications of hierarchical forecasting require practitioners to simplify the two-stage reconciliation process by favoring end-to-end approaches that simultaneously fit all levels of the hierarchy, while still achieving coherence. The end-to-end approach refers to training a model constrained to achieve coherence by directly optimizing for accuracy. End-to-end methods offer advantages such as reduced complexity, improved computational efficiency, and improved adaptability by streamlining the entire forecasting pipeline into a single, unified model. More importantly, end-to-end models generally achieve better accuracy compared to two-stage models that are first trained independently for optimized accuracy and then made coherent through various reconciliation approaches [44, 38].
@@
To our knowledge, only three methods produce coherent probabilistic forecasts and allow models to be trained end-to-end: Rangapuram et al. [44], Olivares et al. [38], and Das et al. [10]. (There is also parallel research on hierarchical forecasts with relaxed constraints [18, 42, 26, 27, 50], but we limit our analysis of this line of work since our focus is on strictly coherent forecasting methods). In particular, Olivares et al. [38] considers a finite mixture of Poisson distributions that captures correlations implicitly through latent variables and does not leverage cross-time series information. Donti et al. [11] achieves coherence by constructing probabilistic forecasts for aggregated time series through the completion of equality. On the other hand, Rangapuram et al. [44] leverages multivariate time series information but achieves coherence through a differentiable projection layer, which could degrade forecast accuracy: it does not directly model correlations between the multivariate outputs, but rather couples them through its projection layer. Dedicated effort is still necessary to capture these hierarchical relationships to improve forecast accuracy. Such probabilistic methods can benefit from the ability to optimize for arbitrary loss functions by differentiating samples, as demonstrated by Rangapuram et al. [44]. The capacity to optimize any loss computed from the forecast samples can help align the forecasting system’s goals with the neural network’s learning objective.
@@
Summarized, an ideal hierarchical forecasting method should satisfy several desiderata: 1) be coherent end-to-end; 2) model a joint multivariate probability distribution, capturing the intricate relationships between series within the hierarchy; 3) leverage cross-time series information and accurately reflect these relationships; and 4) generate differentiable samples, to enable the method to optimize for arbitrary learning objectives that align with the forecasting system’s goals. In this paper, we present a method that satisfies all these ideal properties; see Table 1 for a summary.
@@
We introduce the Coherent Learning Objective Reparameterization Neural Network (CLOVER), a method for producing probabilistic coherent forecasts that satisfies all the desired properties stated above. Our main contributions are the following.
@@
    We design a multivariate Gaussian factor model that achieves forecast coherence exactly by construction, representing the joint forecast distribution over the bottom-level series. This factor model is generic and can augment most neural forecasting models with minimal modifications, with the condition that all the hierarchy’s series are available during training and inference.
@@
    In contrast to other joint distribution models, our factor model samples’ differentiability—enabled by the reparameterization trick—offers greater versatility, allowing optimization of a wider range of learning objectives beyond negative log likelihood. We leverage this versatility by training the model using both the Continuous Ranked Probability Score (CRPS) and the Energy Score [33]. For hierarchical forecasting tasks, aligning the training objective with the CRPS evaluation metric leads to significant improvements in accuracy.
@@
    We combine the multivariate factor model with an augmented MQCNN architecture [53, 37] into the Coherent Learning Objective Reparameterization Neural Network (CLOVER). Our augmented architecture accepts vector autoregressive (VAR) inputs through a multi layer perceptron module.
@@
    CLOVER achieves state-of-the-art performance in six publicly available benchmark datasets, demonstrating substantial improvement in both probabilistic and mean forecast accuracy. On average, CLOVER outperforms the second-best method by 15 percent in probabilistic forecasting accuracy and delivers a 30 percent increase in mean forecast accuracy.
@@
The aggregation matrix ${\mathbf{A}}_{[a][b]}$ represents the collection of linear transformations for deriving the aggregates and sums the bottom series to the aggregate levels. The hierarchical aggregation constraints matrix ${\mathbf{S}}_{[i][b]}$ is obtained by stacking ${\mathbf{A}}_{[a][b]}$ and the $N_{b}\times N_{b}$ identity matrix ${\mathbf{I}}_{[b][b]}$.
For a simple example, consider $N_{b}=4$ bottom-series, so $[b]=\{1,2,3,4\}$ and $y_{Total,t}=\sum_{i=1}^{4}y_{i,t}$. Figure 1 shows an example of such hierarchical structure, where the multivariate hierarchical time series is defined as:
@@
The hierarchical forecasting task augments the forecast probability in Eqn. 3 with coherence constraints in  Eqn. 2.1 [7, 41, 37], by restricting the probabilistic forecast space to assign zero probability to non-coherent forecasts. Definition 2.1 formalizes the intuition, stating that the distribution of a given aggregate random variable is exactly the distribution defined as the aggregates of the bottom-series distributions through the summation matrix ${\mathbf{S}}_{[i][b]}$.
@@
Let $(\Omega_{[b]},\mathcal{F}_{[b]},\mathbb{P}_{[b]})$ be a probabilistic forecast space (on the bottom-level series $\mathbf{Y}_{[b]}$), with sample space $\Omega_{[b]}$, event space $\mathcal{F}_{[b]}$, and $\mathbb{P}_{[b]}$ a forecast probability. Let $\mathbf{S}_{[i][b]}(\cdot):\Omega_{[b]}\mapsto\Omega_{[i]}$ be the linear transformation implied by the aggregation constraints matrix. A coherent forecast space $(\Omega_{[i]},\mathcal{F}_{[i]},\mathbb{P}_{[i]}$) satisfies
@@
for any set $\mathcal{B}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{F}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}_{[b]}$ and its image ${\mathbf{S}}_{[i][b]}(\mathcal{B})\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{F}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}_{[i]}$.
@@
In this work and most of the hierarchical forecast literature, the performance of probabilistic forecasts is primarily evaluated by the Continuous Ranked Probability Score (CRPS), e.g. [6, 44, 38, 41, 10, 54]. The CRPS between a target $y$ and distributional forecast $Y$ is defined as
@@
CRPS is commonly used because it is a strictly proper scoring rule [17] and is agnostic to model and distributional assumptions. The metric also gives a summary of performance on all quantile forecasts [30], as
@@ 3 Methodology
In this section, we describe our main method, the Coherent Learning Objective Reparameterization Neural Network (CLOVER). It consists of a multivariate probabilistic model, an underlying neural network structure, and a coherent end-to-end model estimation procedure.
3.1 Coherent Probabilistic Model
Our predicted probabilistic forecasts at all hierarchical levels are jointly represented by a Gaussian factor model. Our neural network maps the known information (past, static and known future) to the location, scale, and shared factor parameters, and the forecasted factor model parameters are designed to model correlations between the bottom-level series, while conditioning on all known information. Our factor model [^2] Early work on factor forecast models augmenting neural networks done by Wang et al. [52] does not ensure coherence. combined with the coherent aggregation in Eqn. 10 directly estimates the multivariate probability of bottom-level series ${\mathbf{y}}_{[b][t+1:t+N_{h}]}$ conditioning on historical, known-future, and static covariates ${\mathbf{x}}^{(h)}_{[b][:t]},\;{\mathbf{x}}^{(f)}_{[b][t+1:t+N_{h}]},{\mathbf{x}}^{(s)}_{[b]}$, i.e.,
@@ 3.1 Coherent Probabilistic Model
After sampling from the multivariate factor, we coherently aggregate the clipped outputs of the network,
@@
Any bottom-level multivariate distribution can be transformed into a coherent distribution through coherent aggregation  [^3] Coherent aggregation can be thought of a special case of bootstrap reconciliation [41] that only relies on a bottom-level forecast distribution..Given a sample $\mathbf{\hat{y}}_{[b]}\sim\mathbb{P}_{[b]}$, a coherent $\mathbb{P}_{[i]}$ distribution can be constructed with the following sample transformation
@@
In other words, it is enough to aggregate the bottom-level forecasts in a bottom-up manner. We include in Appendix A a proof of the approach’s coherence property and the covariance structure.
@@ 3.2 Neural Network Architecture
Figure 2: The Coherent Learning Objective Reparameterization Neural Network is a Sequence-to-Sequence with Context network that uses dilated temporal convolutions as the primary encoder and multilayer perceptron based decoders for the creation of the multi-step forecast. CLOVER coherently aggregates the samples of the factor model $\mathbf{\tilde{y}}_{[i],\eta,t}=\mathbf{S}_{[i][b]}\mathbf{\hat{y}}_{[b],\eta,t}$. We mark in red the standard normal samples that are parameter-free, the reparameterization trick allows to apply backpropagation through the factor model outputs. CLOVER extends upon the univariate MQCNN, through the cross series multi layer perceptron.
As mentioned in Section 1, the factor model can augment most neural forecasting models if all series in the hierarchy are available during training and inference, as this is a sufficient condition to obtain the factor model parameters. Appendix F shows an augmented subset of NeuralForecast models [36], while the main paper focuses on the MQCNN-based model [53, 37]. We select MQCNN for its outstanding performance in multi-step forecasting problems. Our architecture based on MQCNN has a main encoder which consists of a stack of dilated temporal convolutions, and it is applied to historical information for all series. In addition, it uses a global multi layer perceptron (MLP) to encode the static and future information. The encoder at time $t$ is described in Eqn. 13 below, and is applied to each disaggregated series:
@@
We use a residual cross series MLP to capture vector autoregressive relationships in the hierarchy with minimal modifications to the architecture  [^5] Approach also inspired by regional forecast at Amazon’s Supply Chain Optimization Technologies  [47].:
@@
CLOVER uses a two-stage MLP decoder: the first decoder summarizes information in the horizon-agnostic context $\mathbf{c}^{(ag)}_{[b],t}$ and the horizon-specific context $\mathbf{c}^{(sp)}_{[b][h],t}$; and the second stage decoder transforms the contexts into the factor model parameters $\left(\bm{{\hat{\mu}}}_{[b][h],t},\;\bm{{\hat{\sigma}}}_{[b][h],t},\mathbf{{\hat{F}}}_{[b][k][h],t}\right)$. Eqn. 15 describes the operations:
@@
For the last step, the network composes the factor model samples using Eqn. 9 and aggregates them (equivalent to bottom-up reconciliation) in Eqn. 10.